Please help me with this substitution problem.....
c+a=24
5c+8a=1614

Answer by stanbon(75887)

substitution problem.....
c+a=24
5c+8a=1614
----
c = 24-a
---
Substitute for "c" and solve for "a":
5(24-a) + 8a = 1614
120 +3a = 1614
3a = 1494
a = 498
==============
Solve for "c":
c = 24-a
c = 24-498
c = -474
----
Note: I have a feeling the post was not posted correctly.
The method I used is correct. Maybe you can correct
the arithmetic.
Cheers,
Stan H.
==================